OpenStudy (anonymous):

A theater has 20 rows of chairs. The first row has 16 chairs, the second has 18, the third has 20, and so on. In total, the theater seats _____ people.

OpenStudy (phi):

if all the rows had 16 chairs there would be 20*16

OpenStudy (phi):

the extra seats are 2+4+...+ 38= 2(1+2+3...+19)

OpenStudy (phi):

sum of numbers 1+2+..+n = n(n+1)/2

OpenStudy (anonymous):

what is the answer

OpenStudy (phi):

do you know how to add?

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

700